OpenStudy (anonymous):

Every ten years, the Bureau of the Census counts the number of people living in the United States. In 1790, the population of the U.S. was 3.93 million. By 1800, this number had grown to 5.31 million. Write an exponential function that could be used to model the U.S. population y in millions for 1790 to 1800. Write the equation in terms of x, the number of decades since 1790.

OpenStudy (ankit042):

y = ke^x +c now we know c = 3.93 y=5.31 for x =10 find value of k

OpenStudy (ybarrap):

\[y(t)=Kexp ^{\lambda t}\] \[y(0)=K=3.93\] \[y(1)=3.93\exp ^{\lambda (1)}=5.31\] From this you can get \(\lambda\) and you're done.

OpenStudy (anonymous):

Do I just solve normally?

OpenStudy (ankit042):

sorry I think @ybarrap method is correct. I made a mistake

OpenStudy (anonymous):

Help? What do I do with the last part?

OpenStudy (ybarrap):

yes, solve for the missing variable, \( \lambda \). Do you know how?

OpenStudy (anonymous):

No... :(

OpenStudy (anonymous):

I don't really know what that little wish bone signifies.

OpenStudy (ybarrap):

\[\exp ^{\lambda (1)}=\frac{ 5.31 }{ 3.93 }\] \[\lambda=\ln(\frac{ 5.31 }{ 3.93 })\]

OpenStudy (ybarrap):

\( \lambda \) is the average growth rate over 10 years.

OpenStudy (ybarrap):

In millions of people, per decade.

OpenStudy (anonymous):

oooohhh I see

OpenStudy (mathstudent55):

@ybarrap Isn't the data point Population = 5,310,000 at year 1800 a data point at x = 10? Why do you have \( \lambda(1) \)?

OpenStudy (ybarrap):

Because the question asked about " Write the equation in terms of x, the number of decades since 1790." So the "1" is the 1st decade since 1790.

OpenStudy (mathstudent55):

Good point. Thanks.

OpenStudy (ybarrap):

np